# The Semantic Hub Hypothesis: Language Models Share Semantic Representations Across Languages and Modalities

### v1

@@ 7 Conclusion
This work proposed and investigated the semantic hub hypothesis, which posits that language models represent semantically similar inputs from distinct modalities near one another in their intermediate layers. We find evidence of this phenomena across multiple language models and data types, and further observe that intervening in this space through the model’s dominant language (usually English) leads to predictable model behavior changes.
@@ Acknowledgments
We thank Alex Gu, Alexis Ross, Alisa Liu, Aryaman Arora, Asma Ghandeharioun, Cedegao E. Zhang, Han Guo, Jack Merullo, Linlu Qiu, Mor Geva, Naman Jain, Ruochen Zhang, Sarah Wiegreffe, Shushan Arakelyan, and Yung-Sung Chuang for discussions and help at various stages of this project. Figure 1 uses icons from flaticon.com.
@@ B.1 Multilingual
We automatically evaluate the sentiment of the generation using a DistillBERT-based (Sanh et al., 2020) model finetuned for multilingual sentiment analysis, [^12] https://huggingface.co/lxyuan/distilbert-base-multilingual-cased-sentiments-student, judge the generation fluency by taking the conditional perplexity of the generation given the prefix from Llama-3.1-70B (Llama-3-Team, 2024), [^13] We also tried using Mistral-Nemo-Base (https://huggingface.co/mistralai/Mistral-Nemo-Base-2407), and found similar trends. and compute the relevance of the generation with the prefix by computing the cosine similarity between the generation and the prefix using a XLM-R-Large-based (Conneau et al., 2020) model finetuned for sentence representation [^14] https://huggingface.co/intfloat/multilingual-e5-large (Wang et al., 2024). All these models support both Spanish and Chinese.
@@ Appendix C Token-Level Multimodal Experiments Using the Logit Lens
We consider the alignment of image tokens (corresponding to patches in the original image) to their description words. First, we use the nouns [^15] Words with NOUN or PROPN tags given by SpaCy’s en_core_web_trf model (Honnibal & Montani, 2017). in the image caption as a coarse description—although they do not precisely describe all patches, they should better align with the patches on average than nouns irrelevant to the image. We consider the same 1000 image captions as in §3.5. For each patch, we compute a patch-caption alignment score by summing over the logit lens probabilities for all the nouns in the caption. We average this alignment score over all patches in all images, separately for each transformer layer. For the irrelevant nouns baseline, for each image, we compute the alignment score with an unrelated caption that has the smallest noun overlap with the groundtruth matching caption (we also normalize the number of nouns so that the score is comparable). Figures 13(a) and 13(b) show that the matching caption better aligns with the image patch representations than an unmatched caption for both LlaVa and Chameleon.
Image segmentation labels, which annotate objects in specific image locations, provide a finer-grained patch description. The setup is similar to captions, but the alignment score is not computed using the correspondence between each patch and each noun in the caption, but each patch with the corresponding object label. We say a label correspond to the patch if more than half of the patch’s pixels have that label. For the irrelevant token baseline, we compute the alignment by aligning each patch with a different randomly chosen object category from all categories. Figures 13(c) and 13(d) show that, for LLaVa, the patches are much better aligned to the corresponding labels than randomly assigned labels (which have near-0 logit lens probability). For Chameleon, this is the case for only one middle layer, and not in a statistically significant way, though, as we showed in §4, Chameleon’s latent space can be reliably steered using English tokens.

### v3

@@ 7 Conclusion
This work proposed and investigated the semantic hub hypothesis, which posits that language models represent semantically similar inputs from distinct modalities near one another in their intermediate layers. We find evidence of this phenomenon across multiple language models and data types, and further observe that intervening in this space through the model’s dominant language (usually English) leads to predictable model behavior changes.
@@ Acknowledgments
We thank Alex Gu, Alexis Ross, Alisa Liu, Aryaman Arora, Asma Ghandeharioun, Cedegao E. Zhang, Freda Shi, Han Guo, Jack Merullo, Jonathan May, Linlu Qiu, Mor Geva, Naman Jain, Ruochen Zhang, Sarah Wiegreffe, Shushan Arakelyan, and Yung-Sung Chuang for discussions and help at various stages of this project. Figure 1 uses icons from flaticon.com. This study was supported by funds from MIT-IBM Watson AI Lab.
@@ B.1 Multilingual
We automatically evaluate the sentiment of the generation using a DistillBERT-based (Sanh et al., 2020) model finetuned for multilingual sentiment analysis, [^13] https://huggingface.co/lxyuan/distilbert-base-multilingual-cased-sentiments-student judge the generation fluency by taking the conditional perplexity of the generation given the prefix from Llama-3.1-70B (Llama-3-Team, 2024), [^14] We also tried using Mistral-Nemo-Base (https://huggingface.co/mistralai/Mistral-Nemo-Base-2407), and found similar trends. and compute the relevance of the generation with the prefix by computing the cosine similarity between the generation and the prefix using a XLM-R-Large-based (Conneau et al., 2020) model finetuned for sentence representation [^15] https://huggingface.co/intfloat/multilingual-e5-large (Wang et al., 2024). All these models support both Spanish and Chinese.
@@ Appendix C Token-Level Multimodal Experiments Using the Logit Lens
We consider the alignment of image tokens (corresponding to patches in the original image) to their description words. First, we use the nouns [^16] Words with NOUN or PROPN tags given by SpaCy’s en_core_web_trf model (Honnibal & Montani, 2017). in the image caption as a coarse description—although they do not precisely describe all patches, they should better align with the patches on average than nouns irrelevant to the image. We consider the same 1000 image captions as in §3.5. For each patch, we compute a scalar patch-caption alignment score (for each layer separately) by summing over the logit lens probabilities for all the nouns in the caption at that image patch position. We average this alignment score over all patches in all images, separately for each transformer layer. For the irrelevant nouns baseline, for each image, we compute the alignment score with an unrelated caption that has the smallest noun overlap with the groundtruth matching caption (we also normalize the number of nouns so that the score is comparable). Figures 13(a) and 13(b) show that the matching caption better aligns with the image patch representations than an unmatched caption for both LLaVA and Chameleon.
Image segmentation labels, which annotate objects in specific image locations, provide a finer-grained patch description. The setup is near-identical to captions, but the scalar alignment score for a given patch is not computed using its logit lens probability of caption nouns, but of its corresponding object label. We say a label correspond to the patch if more than half of the patch’s pixels have that label. For the irrelevant token baseline, we compute the alignment by aligning each patch with a different randomly chosen object category from all categories. Figures 13(c) and 13(d) show that, for LLaVA, the patches are much better aligned to the corresponding labels than randomly assigned labels (which have near-0 logit lens probability). For Chameleon, this is the case for only one middle layer, and not in a statistically significant way, though, as we showed in §4, Chameleon’s latent space can be reliably steered using English tokens.